# Sub-shot noise sensitivity via deformed four-headed kitten states

### v2

@@ II.2 Sub-Planckian sensitivity
We now include the example of the Zurek compass state [13], which is recognized as a superposition of coherent states given by $\alpha_{1}=\nicefrac{{c_{0}}}{{\sqrt{2}}}$, $\alpha_{2}=\nicefrac{{-c_{0}}}{{\sqrt{2}}}$, $\alpha_{3}=\nicefrac{{\text{i}c_{0}}}{{\sqrt{2}}}$, and $\alpha_{4}=\nicefrac{{-\text{i}c_{0}}}{{\sqrt{2}}}$, with $c_{0}\in\mathbb{R}^{+}$. This superposition can also be interpreted as the superposition of two cat states, or equivalently as the superposition of four coherent states and is denoted as
@@
with
$G_{k,l}:=\exp\bigg{[}-\frac{1}{2}\left\{|k|^{2}+|l|^{2}\right\}\bigg{]}$ (7)
@@
$W_{\ket{\diamondsuit}}(\beta)=\frac{1}{N_{\diamondsuit}}\sum^{4}_{i,j=1}W_{\ket{\alpha_{i}}\bra{\alpha_{j}}}(\beta)$ (8)
with
$-\alpha_{i}\beta^{*}\big{)}\big{]}.$ (9)
We present $W_{\ket{\diamondsuit}}(\beta)$ in Fig. 1 for a few $c_{0}$ values. As depicted in Fig. 1(a), for $c_{0}=1$, the four coherent states in the compass state cannot be distinguished individually; hence, the corresponding Wigner distribution shows a central positive peak. This scenario represents the four-component kitten state, or alternatively may also referred as a four-headed kitten state [66]. However, for $c_{0}=5$, these four coherent states are now well separated and appear as four Gaussian lobes in the phase space, and the interference pattern is now pronounced in the phase space, as shown in Fig. (1)(b). Figure 1(c) shows that increasing the macroscopic parameter to $c_{0}=8$ causes coherent states to be pushed further away from the phase-space origin, resulting in enhanced negative regions in the intensity plot. Figure 2 illustrates the central interference pattern of each case presented in Fig. 1, where the phase-space features are arranged in a tiled format. Each tile in this pattern have an extension considerably smaller than that of a coherent state, indicating that these structures are at the sub-Planck scale. Furthermore, as $c_{0}$ increases, the size of the sub-Planck features decreases.
@@
The enhancement in sensitivity for a cat state is limited to a particular direction in the phase space. Consequently, compass states are deemed more advantageous than cat states as they may offer greater sensitivity [23]. The overlap between the compass state $\ket{\diamondsuit}$ and its displaced version $\hat{D}(\delta)\ket{\diamondsuit}$ reflects the sensitivity to displacement in phase space,
$S_{\diamondsuit}(\delta)=\left|\sum^{4}_{i,j=1}O_{\ket{\alpha_{i}}\bra{\alpha_{j}}}(\delta)\right|^{2},$ (10)
where
Note that we skip normalization for the sensitivity cases as their intensity visualizations are already normalized to unity. Figures 3(a) and (b) show that the overlap function $S_{\diamondsuit}(\delta)$ is zero for $|\delta|<1$, with $c_{0}=5$ and $c_{0}=8$ in the respective cases involving the sub-Planck structures. This means that the sensitivity to displacement for this compass state is enhanced as compared to a coherent state. Furthermore, compared to the coherent state, the overlap function is now dependent on $c_{0}$. Raising this parameter causes the overlap function to be zero for smaller values of $|\delta|$, indicating increased sensitivity to displacement compared to coherent states. Generally, the sensitivity of a coherent state |$\ket{\alpha}$ is independent of the specific value of $\alpha$, which relates to the average photon number by $|\alpha|^{2}$. Consequently, increasing the average photon number does not enhance the sensitivity of a coherent state to displacements, which is solely limited by the shot noise introduced by vacuum fluctuations [17].
@@
In summary, for the compass state under consideration, sensitivity to displacement appears to be associated with the macroscopic parameter $c_{0}$; as $c_{0}$ grows, it increases both sensitivity and the average photon number in the state. This suggests that a compass state with a larger average photon number can have greater susceptibility to displacement. Furthermore, in this example, the sensitivity to displacement is anisotropically amplified, as illustrated by the tile-like structures around the origin in Fig. 3. Our main focus is on the four-headed kitten state, which is shown in Fig. 1(a). It is evidenced that this particular case is devoid of negative amplitudes, and the dimensions of its phase-space feature are also comparable to those of the coherent states given that sub-Planck structures are absent in this case, and our multiphoton illustrations are specifically devoted to this case, and we will demonstrate how these photon operations will alter its phase-space characteristics by keeping $c_{0}$ a constant.
@@
Figure 4: The overlap between the coherent and deformed versions: (a) For the SA version of the coherent state, the red dotted line represents $r=4$, $q=0$, and the blue solid line represents $r=4$, $q=4$; (b) For the AS version of the coherent state, the red dotted line corresponds to $r=4$, $q=0$, and the blue dotted line represents $r=4$, $q=4$. The horizontal black solid lines indicate the case where $r=0$ and $q=0$, representing the overlap between two ordinary coherent states.
@@ III Refined phase-space densities
Multiphoton operations have been used both theoritically and experimentally to enhance the phase-space characteristics of a quantum state. Techniques involving photon addition (or subtraction) to squeezed-vacuum states have been effectively utilized to generate Schrödinger cat states [50, 51, 52, 53] and can also produce multi-component cat states [48, 36], highlighting the benefits of multi-photon operations. In our current work, we use photon addition and subtraction operations on quantum states to create innovative nonclassical quantum states. These newly proposed states could potentially align with modern experimental setups, as discussed in [52, 53]. In the following sections, we present the quantum states that are the focus of our study, accompanied by a thorough theoretical investigation. This detailed analysis explores the phase-space characteristics and implications of these quantum states, providing a comprehensive understanding of their characteristics and significance.
Figure 5: (a)-(c) represents Wigner distributions of the SA case of a coherent state, and (d)-(f) represents the corresponding AS cases: (a) $r=0$, $q=0$; (b) $r=4$, $q=2$; (c) $r=4$, $q=4$; (d) $r=0$, $q=1$; (e) $r=4$, $q=2$; and (f) $r=4$, $q=4$. In all cases $\alpha=\nicefrac{{1}}{{\sqrt{2}}}$.
III.1 Deformed coherent states
In this section, we review the basic concepts that lead to our primary quantum states; especially, we discuss the concept of deformed coherent states [54, 55, 56]. Deformed coherent states can be generated by applying a sequence of photon addition and subtraction, or vice versa, to a standard coherent state.
Consider the scenario where $r$ photons are first added to a state, followed by the subtraction of $q$ photons. This sequence appears as subtraction and addition (SA) operations and is denoted for a coherent state as
@@ III.1 Deformed coherent states
In the addition-subtraction (AS) scenario, the process involves first subtracting $q$ photons from a state, followed with the addition of $r$ photons. This sequence of operations for a coherent state can be denoted as
|$\ket{\ominus}:=N^{-\nicefrac{{1}}{{2}}}_{\ominus}\hat{a}^{\dagger r}\hat{a}^{q}\ket{\alpha}$ (14)
with
@@
denotes the corresponding normalization coefficient.
To illustrate the impact of photon addition and subtraction operations on a coherent state, we calculate the overlap between the coherent state and its SA and AS variants. The overlap between the coherent state |$\ket{\alpha}$ and its SA version $\ket{\oplus}$ is denoted as $F_{\oplus}(\alpha):=|\braket{\alpha}{\oplus}|^{2}$, while the overlap in the AS case is denoted as $F_{\ominus}(\alpha):=|\braket{\alpha}{\ominus}|^{2}$. These overlaps are evaluated numerically and are represented in Fig. 4. This illustrates that both SA and AS situations differ significantly from the coherent in lower $\alpha$ ranges, and as $\alpha$ grows, corresponding SA and AS variants of the coherent state simply transofrm back to the ordinary coherent state. This comparison is carried out with different amounts of photon addition $r$ and subtraction $q$, as shown in Fig. 4(a) and Fig. 4(b), respectively. Note that the photon subtraction operation in the AS scenario has no effect, and in this case, it essentially represents the photon-added version of the coherent state. This is shown in Fig. 4(b), where the red dotted line and blue dotted line indicate different $q$ values with the same $r$ values, resulting in the same curve.
The Wigner functions of SA and AS cases can be evaluated using Eq. (1). Compared to their counterparts of ordinary coherent states, these Wigner functions attain non-Gaussian form and may hold negative phase-space attributes for different values of $r$ and $q$, highlighting their nonclassical nature [54].
First, for the SA example, the relevant Wigner function is derived as
$W_{\ket{\oplus}}(\beta)=$ $\frac{1}{N_{\oplus}}W_{\ket{\alpha}}(\beta)\sum^{r}_{n=0}\Gamma H_{r-n,q}\left[-\text{i}\Omega,-\text{i}\alpha^{*}\right]$
@@
where
$\Omega=2\beta-\alpha.$ (17)
The non-Gaussian nature of this Wigner function is obvious from the expression, and as it is also evident in Figs. 5(a)-(c) that this Wigner function now acquired negative amplitudes, suggesting the nonclassical nature of this state. Here, note that Fig. 5(a) with $r=0$ and $q=0$, presents the case of corresponding ordinary coherent state, but as observed in Figs. 5(b) and 5(c), when $r$ grows, negative regions are amplified, while the $q$ parameter has the opposite impact. Mathematically, the Wigner function for the AS case has an analogous form to that of the SA case.
$W_{\ket{\ominus}}(\beta)=$ $\frac{(-1)^{r}}{N_{\ominus}}W_{\ket{\alpha}}(\beta)\sum^{r}_{n=0}(-1)^{n}\Gamma(2\beta-\alpha)^{r-l}$
×$\times(\alpha^{*}-2\beta^{*})^{r-l}.$ (18)
This Wigner function is shown in Figs. 5(d)-(f) with different $r$ and $q$ values, indicating that the corresponding Wigner functions also contain negative amplitudes in the phase space. Note that in this case $q$ has no effect on the state, as shown in Fig. (5)(d) with $r=0$ and $q=2$ appearing to have the same Wigner function as a coherent state. This is additional evidence that the photon subtraction effect in the indicated AS case is zero, which simply reflects the photon-added case. The photon subtraction from a coherent state leaves the state uncnhaged is experimentally proved [57]. As illustrated in Fig. (5)(e), similar to the SA case, the photon addition in this case also enhances the negative regions. Furthermore, as observed in Fig. (5)(f), with a nonzero value of $r$, the parameter $q$ remains ineffective. The effects of $r$ and $q$ on the phase space are more thoroughly discussed for the quantum states presented in the following sections.
To emphasize that, as presented in Fig. 5, it is clear that both SA and AS cases presented above are non-Gaussian, and the negativity in their Wigner functions confirms their nonclassical nature, which lacks in the original coherent states. The addition and subtraction (or subtraction and addition) of an equal number of photons from a quantum state can result in two different quantum states. This is confirmed here by two AS and SA cases of coherent states with equivalent photon operations resulting in distinct quantum states, as shown by their Wigner function graphs, and can also be confirmed by the non-commutativity of the bosonic operators $\hat{a}^{\dagger}$ and $\hat{a}$.
Figure 6: PND for SA and AS cases are represented by $P_{\ket{\oplus}}(n)$ and $P_{\ket{\ominus}}(n)$, respectively. (a) $r=0$, $q=0$; (b) $r=4$, $q=2$; (c) $r=4$, $q=4$; (d) $r=0$, $q=2$; (e) $r=4$, $q=2$; and (f) $r=4$, $q=4$. We use $\alpha=\nicefrac{{1}}{{\sqrt{2}}}$ in all cases, and distributions are normalized to unity.

### v3

@@ II.3 Sub-Planckian sensitivity
We now include the example of the Zurek compass state [13], which is recognized as a superposition of coherent states given by $\alpha_{1}=\nicefrac{{c_{0}}}{{\sqrt{2}}}$, $\alpha_{2}=\nicefrac{{-c_{0}}}{{\sqrt{2}}}$, $\alpha_{3}=\nicefrac{{\text{i}c_{0}}}{{\sqrt{2}}}$, and $\alpha_{4}=\nicefrac{{-\text{i}c_{0}}}{{\sqrt{2}}}$, with $c_{0}\in\mathbb{R}^{+}$. This superposition can also be interpreted as the superposition of two cat states, or equivalently as the superposition of four coherent states and is defined as
@@
$W_{\ket{\diamondsuit}}(\beta)=\frac{1}{N_{\diamondsuit}}\sum^{4}_{i,j=1}W_{\ket{\alpha_{i}}\bra{\alpha_{j}}}(\beta).$ (13)
[…]
@@
The overlap between the compass state $\ket{\diamondsuit}$ and its displaced version $\hat{D}(\delta)\ket{\diamondsuit}$ reflects the sensitivity to displacement in phase space
$S_{\diamondsuit}(\delta)=\left|\sum^{4}_{i,j=1}O_{\ket{\alpha_{i}}\bra{\alpha_{j}}}(\delta)\right|^{2}.$ (14)
Figures 3(a) and 3(b) demonstrate that the overlap function $S_{\diamondsuit}(\delta)$ is zero for $|\delta|<1$ along any direction in the phase space, with $c_{0}=5$ and $c_{0}=8$ in the corresponding situations involving the sub-Planck structures. This means that the sensitivity to displacement for this compass state is increased when compared to coherent states as well as when compared to cat states because the corresponding enhancement occurs along all phase-space directions. Furthermore, compared to the coherent state, the overlap function is now dependent on $c_{0}$. Increasing this parameter causes the overlap function to be zero for smaller values of $|\delta|$, which leads to a greater enhancement in the corresponding sensitivity. The enhanced sensitivity achieved by this compass state exhibits anisotropy. This becomes apparent when observing the central patches in Figs. 3(a) and 3(b), where the structures appear tilelike. Further evidence of this anisotropy can be clearly observed in Fig. 12(j) of Appendix A, where the directional dependence is apparent. This indicates that the sensitivity enhancement does not occur uniformly in all directions but rather is more pronounced in certain orientations in phase space.
II.4 Measurement strategies
In this section we mainly elaborate how the detection of weak forces is influenced in a detection scheme that utilizes nonclassical states, with particular emphasis on cat states versus compass states. The precision of quantum parameter estimation is influenced by the energy resources employed during the measurement, such as the average photon number. As indicated in §II.1, the sensitivity of a coherent state |$\ket{\alpha}$ is independent of the specific value of $\alpha$ [Eq. (5)], which is proportional to the average photon number by $|\alpha|^{2}$. This implies that increasing the average photon number is irrelevant to enhancing the sensitivity of a coherent state to displacements, as it is limited by shot noise from vacuum fluctuations [16], a limitation that can be overcome using quantum effects such as superposition and squeezing [32]. For the same average photon number, cat states outperform coherent states. However, compass states show superior sensitivity to displacements compared to both coherent and cat states [as also indicated in Figs. 12(i) and 12(j)], enhancing precision in weak-force detection [28, 29]. In the following, we precisely elaborate on how these detection schemes work.
The concept of weak force detection can be easily understood by considering that a signal one wish to measure is linearly coupled to a harmonic oscillator, causing the oscillator to experience a displacement proportional to the strength of the signal. The sensitivity of the oscillator to these displacements determines the strength of the signal, meaning that quantum states with finer phase-space features (greater sensitivity) are capable of detecting weaker signals. From Eq. (5), it is clear that coherent states can measure displacements of order $|\delta|\sim 1$, setting the standard classical limit for measurement schemes. However, for horizontal cat states with comparatively finer features, as observed in Fig. 12(i), the displacement along the $x$ axis behaves similarly to a coherent state, as in this direction the extension of the patch is also the same as that of the coherent state, while along the momentum axis, the central patch exhibits finer resolution compared to the coherent state, and in this direction the sensitivity is inversely proportional to $c_{0}$, which, for $c_{0}\gg 1$, exceeds the classical limit, also called the Heisenberg limit of the sensitivity. The sensitivity to displacement for compass states scales similarly to that of cat states. However, as shown in Fig. 12(i), unlike cat states, this enhancement is independent of orientation in phase space, allowing compasses to detect weak forces regardless of their direction.
Cat and compass states have been utilized to detect tiny displacements and rotations in phase space, achieving the Heisenberg limit of the sensitivity, especially in cavity systems and systems based on ion traps [96, 97, 29, 27]. For example, in the case of detecting weak forces by using cavity QED and ion trap systems [29, 28], it has been demonstrated that quantum states with coarser phase-space structures, such as cat states, have disadvantages compared to those with finer phase-space features, like sub-Planck structures in such measurement schemes. More specifically, the sensitivity of cat states gradually decreases as the direction of the perturbing force deviates from the axis orthogonal to the line joining the two coherent states.
In metrological applications, achieving simultaneous accuracy in parameters that typically have a trade-off in their measurement precision, such as conjugate variables or other related parameters [98], is a critical challenge. Compasses, on the other hand, have the capacity to get the highest precision in the measurement of conjugate variables at the same level, which has been analyzed in the aforementioned cases  [29, 28]. Additionally, a conventional interferometer can provide an unambiguous estimation of displacement in a known direction in phase space [99, 100, 101]. A hybrid oscillator-qubit interferometric setup using compass state as its local state has demonstrated advantages over the conventional interferometer, as this protocol achieved the unambiguous estimation of phase-space displacements in an unknown direction of a mechanical oscillator [52].
@@ II.4 Measurement strategies
In summary, quantum compasses have proven to be more significant than cat states because of their finer phase-space resolution, which distinguishes them in real-world applications. This distinction highlights the growing importance of quantum states capable of achieving enhanced phase-space resolution, offering distinct advantages in various quantum technologies and measurement protocols. The Zurek compass state maintained sensitivity to displacement [13], which appeared to be associated with the parameter $c_{0}$; as the amplitude $c_{0}$ grows, it increases both sensitivity and the average photon number in the state. This suggests that a compass state with a larger average photon number can have greater susceptibility to displacement [47]. Furthermore, in this example, the sensitivity to displacement is anisotropically amplified, as illustrated by the tilelike structures around the origin in Fig. 3 (see Appendix A for the details). Our main focus is on the case shown in Fig. 1(a), and as it appears, this particular case is devoid of negative amplitudes, and the dimensions of its phase-space feature are also comparable to those of the coherent states given that sub-Planck structures are absent in this case. Our multiphoton illustrations are specifically devoted to this case, and we will demonstrate how these photon operations will alter its phase-space characteristics. The case depicted in Fig. 1(a) can be simply called a four-headed kitten state.
@@
Figure 4: Overlap between the coherent and deformed versions. (a) For the SA version of the coherent state, the red dotted line represents $r=4$ and $q=0$ and the blue solid line represents $r=4$ and $q=4$. (b) For the AS version of the coherent state, the red dotted line corresponds to $r=4$ and $q=0$ and the blue dashed line represents $r=4$ and $q=4$. The horizontal black solid lines indicate the case where $r=0$ and $q=0$, representing the overlap between two ordinary coherent states.
Figure 5: (a)-(c) Wigner distributions of the SA case of a coherent state and (d)-(f) corresponding AS cases: (a) $r=0$ and $q=0$, (b) $r=4$ and $q=2$, (c) $r=4$ and $q=4$, (d) $r=0$ and $q=1$, (e) $r=4$ and $q=2$, and (f) $r=4$ and $q=4$. In all cases $\alpha=\nicefrac{{1}}{{\sqrt{2}}}$.
@@ III Refined phase-space densities
Multiphoton operations have been used both theoritically and experimentally to enhance the phase-space characteristics of a quantum state. For example, techniques involving photon addition (or subtraction) to squeezed-vacuum states have been effectively utilized to generate Schrödinger cat states [73, 74, 75, 76] and can also produce multicomponent cat states [70, 49], highlighting the benefits of multiphoton operations. This section focuses on the quantum states obtained by performing photon addition and subtraction operations in different orders.
Figure 6: PNDs for SA and AS cases of coherent states represented by (a)-(c) $P_{\ket{\oplus}}(n)$ and (d)-(f) $P_{\ket{\ominus}}(n)$, respectively, for (a) $r=0$ and $q=0$, (b) $r=4$ and $q=2$, (c) $r=4$ and $q=4$, (d) $r=0$ and $q=2$, (e) $r=4$ and $q=2$, and (f) $r=4$ and $q=4$. We use $\alpha=\nicefrac{{1}}{{\sqrt{2}}}$ in all cases, and distributions are normalized to unity.
III.1 Degaussification of coherent states
In this section, we discuss the deformed coherent states [77, 78, 79] obtained by applying a sequence of photon addition and then subtraction, or vice versa, to a standard coherent state. First, consider the case when $r$ photons are added to the coherent state, followed by the subtraction of $q$ photons. This sequence appears as subtraction and addition (SA) operations and mathematically, this case for the coherent |$\ket{\alpha}$ is represented as
@@ III.1 Degaussification of coherent states
In the addition-subtraction (AS) scenario, the process involves first subtracting $q$ photons from the state, followed by the addition of $r$ photons, and this case for a coherent state is defined as
|$\ket{\ominus}:=N^{-\nicefrac{{1}}{{2}}}_{\ominus}\hat{a}^{\dagger r}\hat{a}^{q}\ket{\alpha},$ (17)
where
@@
is the corresponding normalization coefficient.
To illustrate the impact of photon addition and subtraction operations on the coherent state |$\ket{\alpha}$, we calculate the overlap between the state |$\ket{\alpha}$ and its SA and AS variants. The overlap between the coherent state |$\ket{\alpha}$ and its SA version $\ket{\oplus}$ is defined as $F_{\oplus}(\alpha):=|\braket{\alpha}{\oplus}|^{2}$, while the overlap in the AS case is defined as $F_{\ominus}(\alpha):=|\braket{\alpha}{\ominus}|^{2}$. These overlaps are evaluated numerically and are represented in Fig. 4, where it is evident that both SA and AS situations differ significantly from the coherent state in lower $\alpha$ ranges, and as $\alpha$ grows, corresponding SA and AS variants of the coherent state simply transform back to the ordinary coherent state. This comparison is carried out with different amounts of photon addition $r$ and subtraction $q$, as shown in Fig. 4(a) and Fig. 4(b), respectively. Note that the photon subtraction operation in the AS case has no effect, and it essentially represents the photon-added version of the coherent state. This is justified in Fig. 4(b), where the red dotted line and blue dashed line correspond to different $q$ but same $r$ values, resulting in the same curve.
The Wigner functions of SA and AS cases are evaluated using Eq. (1). Compared to their counterparts of ordinary coherent states, these Wigner functions attain non-Gaussian form and may exhibit negative phase-space attributes for different values of $r$ and $q$, highlighting their nonclassical nature [77, 102, 103]. First, for the SA example, the corresponding Wigner function is derived as
$W_{\ket{\oplus}}(\beta)=$ $\frac{1}{N_{\oplus}}W_{\ket{\alpha}\bra{\alpha}}(\beta)\sum^{r}_{n=0}\Gamma H_{r-n,q}\left[-\text{i}\Omega,-\text{i}\alpha^{*}\right]$
@@
where $\Omega=2\beta-\alpha$. The non-Gaussian nature of this Wigner function is obvious from the expression (III.1), and it is also evident in Figs. 5(a)-5(c) that this Wigner function has now acquired negative amplitudes, reflecting the nonclassical nature of the SA case of the coherent state. Note that Fig. 5(a) with $r=0$ and $q=0$, represents the case of the corresponding ordinary coherent state, but as observed in Fig. 5(b) for nonzero $r$ and $q$ values, the corresponding Wigner function holds negative amplitudes, and then in Fig. 5(c) an increment in $q$ values with a constant $r$ reduces the negative amplitudes in the Wigner function. For this case, it can be shown that higher $r$ amplifies negative regions while larger $q$ may cause the elimination of the negative amplitudes in the Wigner phase space.
Mathematically, the Wigner function for the AS case has a form analogous to that of the SA case, that is,
$W_{\ket{\ominus}}(\beta)=$ $\frac{(-1)^{r}}{N_{\ominus}}W_{\ket{\alpha}\bra{\alpha}}(\beta)\sum^{r}_{n=0}(-1)^{n}\Gamma(2\beta-\alpha)^{r-n}$
×$\times(\alpha^{*}-2\beta^{*})^{r-n}.$ (20)
This Wigner function is shown in Figs. 5(d)-5(f) with different $r$ and $q$ values, indicating that the corresponding Wigner functions also contain negative amplitudes in the phase space. Note that in this case $q$ has no effect on the state, as shown in Fig. (5)(d), with $r=0$ and $q=1$ appearing to have the same Wigner function as a coherent state. This is additional evidence that the photon subtraction effect in the indicated AS case is zero, and we know that it simply reflects the photon-added case. The photon subtraction from a coherent state leaves the state unchanged, as has been proved experimentally [80], as the coherent state is an eigenstate of the annihilation operator. As illustrated in Fig. (5)(e), similar to the SA case, the photon addition in this case also enhances the negative regions. Furthermore, as illustrated in Fig. (5)(f), where the parameter $r$ is kept the same as in the previous case and $q$ is increased, again it appears that the increase in $q$ (the number of photon subtraction) in the AS case has no effect on the Wigner distribution.
In summary, both SA and AS cases of the coherent state are non-Gaussian, and the negativity in their Wigner functions confirms the nonclassical nature of these states, which is lacking in the original coherent states. The addition and subtraction (or subtraction and addition) of an equal number of photons from a quantum state can result in two different quantum states. This is confirmed here by two AS and SA cases of coherent states with equivalent photon operations resulting in distinct quantum states, as shown by their Wigner function illustrations, and can also be confirmed by the noncommutativity of the bosonic operators $\hat{a}^{\dagger}$ and $\hat{a}$.
Figure 7: Overlap function $|F_{\square}(c_{0})|^{2}$ (blue dashed curve) and $|F_{\blacksquare}(c_{0})|^{2}$ (red solid line), with $r=4$ and $q=4$. The horizontal black dashed line depicts the case where $r=0$ and $q=0$; consequently, the overlap is simply between two compass states.